The value of KSp of CaCO3 and CaC2O4 are 4.7x10-9 and            1.3 x10-9 respectively at 25degree c.If the mixture of these two is washed with water ,what is the concentration of Ca+2 ions in water

Grade:

2 Answers

Naveen Kumar
9 years ago
Dear student also specify the amount of water added.
Suppose we have solubility of these compounds x and y respectively.
CaCO3................>Ca2++ CO32-
…............................x.............x
x*x=4.7*10^-9, x=6.7*10^-5

CaC2O4..................> Ca2++ C2O42-
….................................y.............y
y*y=1.3*10^-9, y=3.6*10^-5
Now if both are mixed, then Ca2+ would be the common.
Naveen Kumar
9 years ago
Dear student also specify the amount of water added.
Suppose we have solubility of these compounds x and y respectively.
CaCO3................>Ca2++ CO32-
…............................x.............x
x*x=4.7*10^-9, x=6.7*10^-5

CaC2O4..................> Ca2++ C2O42-
….................................y.............y
y*y=1.3*10^-9, y=3.6*10^-5
Now if both are mixed, then Ca2+ would be the common.